OpenStudy (anonymous):

question on writing equations: The perimeter of a rectangle is 482 cm. The length is 6 cm greater than 4 times the width. Find the length.

OpenStudy (anonymous):

The perimeter of a rectangle is 482 cm 2*L+2*w = 482 The length is 6 cm greater than 4 times the width. L=4*w+6 Find the length. Solve the following equations for L and w: {2*L+2*w = 482, L=4*w+6} {L -> 194, w -> 47}

OpenStudy (anonymous):

Length = 2(4w +6) Width = 2w Set the equation up. 2w + 2(4w +6) = 482 Find the w (width) 2w + 2(4w +6) = 482 2w + 8w + 12 = 482 10w + 12 = 482 10w = 482-12 10w = 470 \[\frac{10w}{10} = \frac{470}{10} \] w (width) = 47 Now that we have the width we look at where says that the Length is 6 cm Greater than 4 times the width So we use the 47, which is our width. So we are looking for 4w +6 plug the 47 into the w because 47 is our width 4(47) + 6 = 194 So our length is 194